OpenStudy (anonymous):

The perimeter of a swimming pool is 96m. The width is 3m more than half the length of the pool. Find the length and width of the pool. (Perimeter of a rectangle = 2 L + 2 W)

OpenStudy (anonymous):

96 = (2L+3) + ((.5)(2W))

OpenStudy (anonymous):

if this helped please become a fan

OpenStudy (radar):

w=3+L/2, L=L 96=2L + 2(3+L/2), 96=2L + 6 +L, 96 = 3L +6, 3L = 96 - 6 =90 L = 90/3 = 30 W= 3 + 30/2 = 18

OpenStudy (radar):

in meters of course